OpenStudy (anonymous):

Which is the equation of the given line in point-slope form (-4,-1) (1/2,3) y+1=8/9(x+4) y+4=9/8(x+1) y+1=8/7(x+4) y-1=8/9(x-4)

OpenStudy (anonymous):

equation of the line passing through (x1,y1) & (x2,y2) is given by (y-y1)/ (y2-y1)=(x-x1)/(x2-x1)

OpenStudy (kkutie7):

(3+1)/(1/2+4)=slope Y=slope(x+4)-1

OpenStudy (anonymous):

Huh?

OpenStudy (anonymous):

y+1=8/9(x+4)

OpenStudy (kkutie7):

Find the slope but this equation (y2-y1)/(x2-x1) that's what I had above. Then put it in the equation for y. Y=slope(x-x1)+y1 you have x1 and y1 from the points in your question.

OpenStudy (kkutie7):

The answer is written in this form though. Y-y1=slope(x-x1)